### v1

# Secure Information Embedding and Extraction in Forensic 3D Fingerprinting

Abstract
The prevalence of 3D printing poses a significant risk to public safety, as any individual with internet access and a commodity printer is able to produce untraceable firearms, keys, counterfeit products, etc. To aid government authorities in combating these new security threats, several approaches have been taken to tag 3D-prints with identifying information. Known as fingerprints, this information is written into the object using various bit embedding techniques; examples include varying the height of the molten thermoplastic layers, and depositing metallic powder with different magnetic properties. Yet, the practicality of theses techniques in real-world forensic settings is hindered by the adversarial nature of this problem. That is, the 3D-printing process is out of reach of any law enforcement agencies; it is the adversary who controls all aspects of printing and possesses the printed object. To combat these threats, law enforcement agencies can regulate the manufacturing of 3D printers, on which they may enforce a fingerprinting scheme, and collect adversarially tampered remains (e.g., fragments of a broken 3D-printed firearm) during forensic investigation.
Therefore, it is important to devise fingerprinting techniques so that the fingerprint could be extracted even if printing is carried out by the adversary. To this end, we present SIDE (Secure Information Embedding and Extraction), a fingerprinting framework that tackles the adversarial nature of forensic fingerprinting in 3D prints by offering both secure information embedding and secure information extraction. First, with the help of ARM TrustZone, SIDE protects the integrity of the information embedding process such that the 3D-print is guaranteed to be tagged with the correct fingerprint. Second, by devising coding-theoretic techniques, SIDE is break-resilient, i.e., it allows correct fingerprint extraction even if the adversary breaks the 3D-print to at most a certain number of parts.
1 Introduction
3D printing brings about a revolution in consumption and distribution of goods. However, the prevalence of this technology exposes the public to various security threats. Any individual with internet access is able to produce untraceable firearms and counterfeits of IP-protected products (Figure. 1[1, 2, 3, 4]. To aid authorities and law enforcement agencies in combating these threats, fingerprinting 3D printed objects is one potential solution. In a fingerprinting scheme, a 3D-printer embeds uniquely traceable data into the printed object (e.g., timestamp, geoposition, printer ID, etc.), that can later be extracted to trace the perpetrator, e.g., if the object was confiscated or left at a crime scene.
Figure 1: Renderings of (a) a 3D printed lower receiver of an AR-15 rifle. While US law regulates commercial distribution of this receiver, it can be easily manufactured privately using 3D printing technology, resulting in untraceable firearms known as ghost guns, and (b) an IP-protected tabletop wargame miniature that is illegally scanned with a 3D scanner.
Various techniques for embedding bits in 3D printed objects have been proposed in the literature. For instance, embedding tags on the surface of the object by altering layer height [5], inserting cavities inside the object [6], altering the surface of the object by varying printing speed [7], embedding RFID-tags [8, 9, 10] or QR codes [11, 12], inserting acoustic barcodes using surface bumps [13], etc. For a broad introduction to the topic see [14, 15].
An orthogonal line of work in the forensic 3D-printing literature proposes to identify printers from 3D-prints using machine-learning methods, based on manufacturing imperfections of printer components reflected on the print results [16]. However, whether the proposed method provides enough resolution for a large number of 3D printers remains unknown. Moreover, [16] focuses on identifying the 3D printer from the printed object, and does not provide a method for embedding arbitrary bits. These bits, which may represent geoposition or timestamp, are unknown at the time of recording the printer’s imperfections, and are crucial for forensic applications.
The practicality of any such bit embedding technique, however, is hindered by a fundamental obstacle, i.e., the adversary controls the printing. Specifically, the adversary has full control over the entire printing process, and may manipulate the prints after committing a crime. For example, the adversary may bypass the information embedding software installed on the printer control board and print a fingerprint-free firearm. The firearm may also be broken after committing the crime, making the embedded information unrecoverable.
In contrast, law enforcement agencies are restricted to merely regulate the manufacturing and distribution of commercial 3D printers. Once the printer is handed over to the (potential) adversary, the agencies can hardly monitor its usage, and can only attempt to recover information from adversarially tampered 3D-prints during forensic crime investigation. Therefore, a practical information embedding scheme must guarantee successful embedding and successful extraction of a reasonable number of bits, even in circumstances where printing is out of reach of law enforcement authorities, and even in circumstances where the 3D-prints are adversarially damaged (Figure 2).
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionpt3D PrinterptLaw Enforcement AgencypttamperptprintptmanufactureptCompaniesptAdversary
Figure 2: Law enforcement agencies only regulate 3D printer manufacturers, and confiscate adversarially tampered 3D-prints. The printing process and the printed object are fully controlled by the adversary, and are beyond the reach of law enforcement.
Indeed, any embedded bits can be stripped-off from the 3D model file before printing, and any tag can be stripped-off using household items, as the printed object itself could be melted, deformed, or simply broken apart. As such, the aforementioned methods fail in the face of a knowledgeable adversary that tampers with the printing process and the prints, as they rely on the integrity of both the information embedding process and the geometric shape of the printed object.
To this end, we propose SIDE, a secure information embedding and extraction framework that is robust against a knowledgeable adversary that controls the printing process and may tamper with the printed object. Specifically, our contribution in this work is two-fold:
  • Secure information embedding: SIDE incorporates a Trusted Execution Environment (TEE) into the 3D printing process in order to protect its integrity. Consequently, the information embedding procedure is guaranteed to succeed even if the adversary is able to compromises the printer.
  • Secure information extraction: SIDE employs coding-theoretic techniques for correct information extraction from the 3D-prints even if their geometric shapes are adversarially damaged. As a result, the law enforcement agency is guaranteed to extract information correctly.
2 Background and Related Work
2.1 3D Printing
3D Printing, commonly referred to as Additive Manufacturing (AM), has emerged as a revolutionary technology with profound implications across various industries. In contrast to the traditional subtractive manufacturing during which materials are consecutively removed from the workpiece, 3D printing refers to an additive process of creating a physical object and is typically done by laying down many thin layers of material (such as liquid plastic or powdered metal) in succession. Indeed, the manufacturing process starts with “slicing” the to-be-manufactured 3D model into discrete horizontal layers/slices using a software called slicer. Each layer represents a planar cross-section of the model along the printing direction (the $z$-direction).
Numerous technologies have been developed for 3D printing. By and large, they differ by which material is in use and how layers are formed. A layer can either be formed by using a nozzle that deposits molten thermoplastics while shifting back and forth on a surface, e.g. Fused Deposition Modeling (FDM), by depositing a layer of liquid polymers and curing it by light exposure, e.g., Stereolithography (SLA), or by depositing powdered material over the surface and binding it using high-energy laser beams, e.g., Selective Laser Sintering (SLS).
In this research, we focus on the FDM technology for its prevalence in commodity 3D printers (compared with SLS), as well as the profound mechanical strength offered to the printed objects (compared with SLA). In an FDM-based printer, the slicer operates by converting the layers to a G-code, which is a series of commands for the printer’s hardware. Such commands may include the nozzle movement along the $x$$y$, and $z$ axes, the extrusion of material, the nozzle/bed temperature, and other specific aspects based on the printer’s capabilities.
It is worth noting that a G-code command specifies only the expected action of the printer hardware in a relatively high level, while the low-level implementation is not addressed. For example, the command
G1 X98.302 Y96.831 E15.796
merely instructs the printer to move its nozzle to $(98.302,96.831)$ in the $x$-$y$ coordinate and simultaneously extrude $15.796$ millimeter of molten thermoplastics. Completing such an operation requires a series of discrete angular-step rotations of multiple stepper motors, which are directly operated by the printer firmware.
Indeed, the firmware plays a crucial role in bridging the G-code and printer hardware. In spite of commands for heaters and other applicable components, the firmware translates the G-code to signals that control the kinematics of the printer’s actuators, e.g., stepper motors. This translation task is non-trivial, and has a significant impact on print quality. For instance, a sudden jerk of the printer nozzle may lead to uneven deposition of print material, which jeopardizes the final print quality or even fails the print. In contrast, a nozzle movement with smooth velocity change is generally preferred. Therefore, the print quality may be significantly improved with an updated firmware, and the same G-code may be executed on printers with completely different builds.
2.2 Existing 3D Fingerprinting Methods
Several methods for embedding bits in 3D-printed objects have recently been proposed in the literature. These technologies allow the printer to vary either the orientation of the nozzle, the thickness of the layer, or the printing speed. Within reasonable bounds, varying either of those has a marginal effect on the functionality of the object. By varying layer thickness, for example, the printer can embed a 00 by printing a layer that is slightly thinner, and a $1$ by printing a layer that is slightly thicker, than some reference thickness. By varying the orientation of the nozzle, bits can be embedded by the relative orientation of adjacent layers; for example, if two adjacent layers are oriented similarly, the embedded bit is 00, and otherwise it is $1$. Both methods are illustrated in Figure 3.
Figure 3: Two possible methods for embedding bits in a 3D printed object with little to no effect on functionality: (a) Embedding by layer thickness; thicker layers represent $1$ and thinner layers represent 00. (b) Embedding bits using the orientation of adjacent layers; if two adjacent layers are oriented similarly, it is a 00, and if oriented differently, it is a $1$. Both left and right figures contain the bits $0101$.
Similar ideas have been implemented successfully in several recent works. Ref. [5] varies the thickness of each layer across several adjacent layers to create a matrix of bits that is visible to the naked eye on the surface of the object. Parity bits were then added to resolve reliability issues in some cases, and additional noise patterns were discussed, such as orientation issues and sanding. In the method LayerCode [17], variations in color and thickness were used to embed a barcode on the surface of the printed object, that can be retrieved using a smartphone camera.
An orientation-based method has been implemented in [18], where the authors print a reference layer that is circularly grooved by a sequence with low auto-correlation. Data is embedded in all other layers by the respective angle of the layer to the reference layer; this enables encoding with alphabet size larger than two.
Various other creative ideas have been explored. Examples include embedding information-carrying cavities within the object [19, 20], water-marking the 3D-mesh of the surface of the object [21], inserting RFID tags [8], inserting a series of notches which create an acoustic barcode when tapped [13], etc. In the data extraction phase, most existing methods rely on an RGB camera, a 3D scanner, or an ordinary scanner. Future technologies however, such as the ones in Figure 3, might require an industrial CT scanner.
However, we argue that none of these approaches is suitable for forensic applications. First, they all implicitly assume that a mechanism for correct information embedding is in place. Such a mechanism is crucial since in most scenarios the adversary owns the printer and/or the file, and might potentially remove the embedded bits altogether. Second, none of the methods is provably resilient to adversarial tampering; they can be easily breached by a knowledgeable adversary that can scrape the object or break it apart.
A different line of works applies machine-learning based methods to identify the intrinsic fingerprint which results from printer hardware imperfections (PrinTracker [16]). The variations of printer components arising from manufacturing errors are essentially inevitable, and create perceivable differences between objects made from different 3D-printers, even if they are printed from the same 3D model. These differences form a unique printer fingerprint, and can be extracted from the printed objects with commodity scanners.
Unlike the aforementioned techniques, PrinTracker is a fingerprinting solution which considers the adversarial nature of forensic applications, as well as potential attacks. This is different from the information embedding solution proposed in our work. Although information embedding can be used for fingerprinting purpose, it is not limited to it. The proposed solution allows for arbitrary bit embedding, which enables a variety of forensic applications. For example, certain applications may require the embedding of timestamp or geoposition of the printing into the printed object in order to check whether the printer has been misused in unauthorized time and location. Further, it remains to be shown if the extraction, storage, and identification of such imperfections scales well to a large number of printers.
2.3 Coding methods
Similar coding problems of information extraction from fragments have been previously studied in the literature, motivated by applications in DNA storage. In particular, several variations of the so-called torn-paper channel were studied in [22, 23, 24, 25, 26], where [22, 23, 24] focused on a probabilistic error model which is incompatible with our adversarial setting, and [25] studied an adversarial error model in which fragment are restricted in length. Most closely related, [26] studied $t$-break codes, which coincide with our problem definition (given shortly). They provided a theoretical analysis of the fundamental limits, and an (almost) matching code construction. However, the scheme described in [26] involves random encoding and is only effective for a very large number of embedded bits. Our work herein addresses the same model of [26], yet at a much simplified framework which can be easily employed in practical settings, albeit below the maximum theoretical information rate.
3 Threat Model
We consider an knowledgeable adversary who wishes to commit a crime. The adversary is unable to obtain a certain required instrument, say, fails to pass background checks when attempting to purchase a firearm, or has insufficient financial means. Hence, the adversary manufactures the instruments independently, by downloading a 3D model file and feeding it to a commercial grade 3D-printer bought from a local electronics retailer.
We assume that the 3D printer, which does not require a background check to purchase, holds a binary string $\mathbf{w}$. The content of such string may include a unique printer ID, the time and geoposition of the printing process, the identity of the file, etc, which aid the authorities to trace the perpetrator111 In the upcoming experiments we focus solely on printer ID. The secure procurement and embedding of additional real-time identifying information is left for future studies. .
The string $\mathbf{w}$ is embedded in every object printed by that printer, in an arbitrary fashion [^2] Our methods are not inherently restricted to any particular bit embedding technique, as long as those bits are manifested as physical elements on the printing direction along the object. For the sake of experiments, we embed bits by varying layer heights.. We assume that the actuators (e.g., stepper motors) and other electronics (e.g., heater) of the 3D-printer are connected to a hardware platform (e.g., a control board) which supports a trusted execution environment (TEE). The TEE provides strong isolation on processor, memory, and peripherals (necessary for connecting the components). The connection from the board to the electronic components is also free of malicious tampering.
Meanwhile, the adversary is able to compromise the rich execution environment (REE) on the control board, including all necessary software components and the hosting operation system. Then, the adversary prints the object, commits the crime, and mechanically damages the printed tool, breaking it into at most $t$ fragments along the printing direction (i.e., the “$z$” direction). The reasoning behind this assumption stems from the following observations.
  1. 1.
    3D printers operate in a layer-by-layer fashion, and the 3D-prints have the lowest mechanical strength along the printing direction. If the break crosses multiple layers, then the irregular geometry of the broken section is likely to enable correct reassembly.
  2. 2.
    The parameter $t$ reflects time constraints, access to instruments, or simply physical strength limitations, which prevent the adversary from breaking the object to arbitrarily many fragments. Hence, the parameter $t$ is expected to be small with respect to the number of embedded bits.
We assume the fragments of the broken instrument are confiscated by a law enforcement agency after the crime has been committed. The agency is able to read the embedded bits from the fragments by examining the height of each layer. Examples of equipment for this task includes microscopes and computed tomography (CT) devices. However, since the bit strings are extracted from at most $(t+1)$ fragments, the agency is not aware of their correct order.
4 SIDE Overview
SIDE incorporates coding-theoretic techniques with hardware isolation mechanism for secure information embedding and secure information extraction.
The resemblance between G-code and assembly code motivates the use of TEE to protect the integrity of the printing process. Indeed, both G-code and assembly code are instructions for lower-level hardware, but both are oblivious to the inner workings of that hardware. For example, an instruction to fetch memory data to a register does not specify how this operation is implemented in the integrated circuit (IC). Meanwhile, two ICs that support the same instruction set may have completely different circuit designs.
For secure extraction of information in case of breaks, the information $\mathbf{w}$ goes through an encoding process before being embedded. The resulting codeword $\mathbf{c}$ has the $t$-break-resilient property; as the name suggests, the corresponding decoder is able to recover $\mathbf{w}$ even if the codeword is sliced into at most $(t+1)$ unordered fragments.
Next, the SIDE slicer generates G-code with an embedded break-resilient codeword using some method (e.g., varying layer height), and the SIDE firmware executes the G-code, i.e., translates it to signals that directly operate the printer’s actuators. For secure embedding purposes, both the generation and the execution of the G-code are isolated in a trusted environment. Then, the result is forwarded to the peripherals that communicate with printer actuators; this process is also assumed to be protected by TEE.
5 Secure Information Extraction
5.1 Overview
The secure information extraction feature of SIDE is attributed to a break-resilient code (BRC), which is specifically developed for forensic fingerprinting purposes. BRC includes an encoder which holds a binary string $\mathbf{w}\in{\{0,1\}}^{k}$ (an information word, e.g., printer ID) for some integer $k$, that is to be encoded to a string $\mathbf{c}\in{\{0,1\}}^{n}$ (a codeword). After encoding $\mathbf{w}$, the codeword $\mathbf{c}$ is embedded in a printed object. BRC may be applied to any possible bit embedding method, and yet in Section 6.2 we focus on varying layer width. For a security parameter $t$, an adversary breaks these $n$ bits at arbitrary $t$ locations or less, resulting in a multiset of at most $t+1$ fragments.
Example 1.
For a string $\mathbf{c}=0100011100$ and $t=3$, the possible fragment multisets include
[…]
These bits are extracted from the fragments of the broken printed object, and then given to the decoder in an unordered fashion. The goal of the decoder is to reconstruct $\mathbf{c}$ exactly in all possible cases, from which $\mathbf{w}$ can be obtained, hereafter providing secure information extraction. The following notions from coding theory are employed as basic building blocks for BRC.
5.1.1 Systematic Reed-Solomon Codes
Systematic Reed-Solomon (RS) codes are a special type of Reed-Solomon codes, which have been widely employed in communication systems and data storage applications; for an introduction to the topic see [27, Ch. 5].
For integers $k$ and $n$ such that $n>k>0$, a systematic $[n,k]$ RS code is a set vectors of length $n$ called codewords, each entry of which is taken from  $\mathbb{F}_{q}$, a finite field with $q$ elements. The first $k$ entries of each codeword in a systematic RS code carry information in raw form, and the remaining $n-k$ contain redundant field elements that are used for error correction. This code is maximum distance separable (MDS), a property which allows the recovering of a codeword, after being corrupted by $a$ errors (incorrect symbols with unknown locations) and $b$ erasures (incorrect symbols with known locations), as long as $n-k\geq 2a+b$.
Note that RS codes requires the finite field size $q$ to be greater or equal to the length of the codeword, i.e., $q\geq n$. Furthermore, we focus on the binary field and its extension field in this paper, in which $q=2^{x}$ for some integer $x\geq 1$, and any element can be represented by a binary string of length $x$.
5.1.2 Run-length Limited Codes
A run-length limited (RLL) code is a collection of codewords in which the length of runs of repeated bits is bounded. Specifically, in this paper we employ the RLL code from [28, Algorithm 1] for its simplicity of implementation.
5.1.3 Mutually Uncorrelated Codes
A mutually uncorrelated (MU) code has the property that for every two (possibly identical) codewords, the prefix of one is not identical to the suffix of another. As such, the codewords of a MU code do not overlap with each other when appearing as substrings of a binary string. MU codes has been extensively investigated in the past [29, 30, 31, 32, 33, 34, 35, 36, 37, 28].
We adopt a classic construction of MU code, in which each codeword starts with ${\lceil\log k\rceil}$ zeros followed by a one, where $k$ is the length of the (binary) information word. The last bit is fixed to one, and the remaining bits are free from zero runs of length ${\lceil\log k\rceil}+1$.
5.1.4 Distinct Strings
Inspired by ideas from [28, Algorithm 1] and [38, Algorithm 1], we provide Algorithm 3 that maps an input word to an array of pairwise distinct binary strings, and offer the inverse operation in Algorithm 4. These procedures serve as an important component in BRC. Both algorithms, as well as proofs of their correctness, are given in the appendix.
Algorithm 1 Encode
1:Input: An information word $\mathbf{w}\in{\{0,1\}}^{k}$, where $k={(l-t)\cdot m-1},$ and $l,m$ are positive integers such that $m\geq{\lceil 2\log l\rceil}+2$.
2:Output: A codeword $\mathbf{c}\in{\{0,1\}}^{n}$, where $n=l\cdot(m+{\lceil\log m\rceil}+4)+t\cdot(4m+11)$
3:Let $\mathbf{u}\leftarrow\mathbf{p}^{(0)}\circ\cdots\circ\mathbf{p}^{(t-1)}\circ\mathbf{w}$, where $\mathbf{p}^{(i)}$ is the binary representation of $i\in[0,t-1]$ in $m$ bits.
4:$\texttt{dStrings}=(\mathbf{u}_{1},\cdots,\mathbf{u}_{l})\leftarrow\textsc{d-encode}(\mathbf{u})$
5:Let next be a key-value store with keys and values being element in ${\{0,1\}}^{m}$.
6:for all keys $\mathbf{s}$ in next in ascending order do
7:     if there exist $i\in[0,l-2]$ such that $\mathbf{s}=\mathbf{u}_{i}$ then
8:          $\texttt{next}[\mathbf{s}]\leftarrow\mathbf{u}_{i+1}$
9:     else
10:          $\texttt{next}[\mathbf{s}]\leftarrow\mathbf{s}$
11:$\mathbf{r}_{1},\cdots,\mathbf{r}_{4t}\leftarrow\textsc{rs-encode}((\texttt{next}[\mathbf{s}_{0}],\cdots,\texttt{next}[\mathbf{s}_{2^{m-1}}]),4t)$, where $\mathbf{s}_{i}=\mathbf{p}^{(i)}.$
12:$\mathbf{c}=\textsc{mu}(\mathbf{u}_{0})\circ\textsc{rll}(\mathbf{d}_{0})\circ\cdots\circ\textsc{mu}(\mathbf{u}_{t-1})\circ\textsc{rll}(\mathbf{d}_{t-1})\circ\textsc{mu}(\mathbf{u}_{t})\circ\cdots\circ\textsc{mu}(\mathbf{u}_{l-1}).$
13:Output $\mathbf{c}$
5.2 Encoding
The encoding procedure of BRC is provided in Algorithm 1. Let $\mathbf{w}\in{\{0,1\}}^{k}$ be an information word, where
$k={(l-t)\cdot m-1},$
and where $l$ and $m$ are positive integers such that $m\geq{\lceil 2\log l\rceil}+2$. The encoding of $\mathbf{w}$ begins by prepending it with $\mathbf{p}^{(0)},\mathbf{p}^{(1)},\cdots,\mathbf{p}^{(t-1)}$, resulting in
$\mathbf{u}=\mathbf{p}^{(0)}\circ\cdots\circ\mathbf{p}^{(t-1)}\circ\mathbf{w}\in{\{0,1\}}^{l\cdot m-1},$ (1)
where $\mathbf{p}^{(i)}$ is the binary representation of $i\in[0,t-1]$ in $m$ bits (line 3). Then, as shown in line 4, the resulting string $\mathbf{u}$ is fed into the function d-encode (which is implemented in Algorithm 3 in the appendix) and mapped to an array of $l$ pairwise distinct binary strings of length $m$, i.e.,
$\texttt{dStrings}=(\mathbf{u}_{0},\cdots,\mathbf{u}_{l-1}).$ (2)
Note that due to the implementation of d-encode, the first $t$ elements of dStrings remain intact, i.e., $\mathbf{u}_{i}=\mathbf{p}^{(0)}$ for all $i<t$, and they are referred as markers. In the next step, a key-value store next is defined to represent the ordering of elements in dStrings as follows (line 610). For every key $\mathbf{s}\in{\{0,1\}}^{m}$, the value $\texttt{next}[\mathbf{s}]$ is defined as
$\texttt{next}[\mathbf{s}]=\begin{cases}\mathbf{u}_{i+1}\leavevmode\nobreak\ &\text{if}\leavevmode\nobreak\ \mathbf{s}=\mathbf{u}_{i}\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ i\in[0,l-2],\\ \mathbf{s}\leavevmode\nobreak\ &\text{otherwise.}\end{cases}$
Note that $\texttt{next}(\mathbf{s})$ is well defined, since $\mathbf{u}_{i}\neq\mathbf{u}_{j}$ for every $i\neq j$ by the pairwise distinct property of dStrings. It is also worth noting that the mapping from dStrings to next is injective; one may recover dStrings from next by observing every key $\mathbf{r}$ such that $\texttt{next}(\mathbf{r})\neq\mathbf{r}$, and connecting every two $\mathbf{r}_{a},\mathbf{r}_{b}$ of them if $\texttt{next}(\mathbf{r}_{a})=\mathbf{r}_{b}$.
We proceed to the treatment of next. Since the values in next are binary strings of length $m+1$, they can be regarded as symbols in the finite field $\mathbb{F}_{2^{m+1}}$. They are sorted by the their corresponding keys and fed into a systematic Reed-Solomon encoder, which then generates $4t$ redundancy strings $\mathbf{r}_{1},\cdots,\mathbf{r}_{t}\in{\{0,1\}}^{m+1}$ (line 11). Note that such encoding is feasible since the codeword length $2^{m}+4t$ is smaller than the number of elements in $\mathbb{F}_{2^{m+1}}$.
The codeword $\mathbf{c}$ consists of two parts. The first region is called the information region, as it is generated from $\mathbf{u}_{0},\cdots,\mathbf{u}_{l-1}$, which directly originate from the information word $\mathbf{w}$. The second region is called the redundancy region. As the name suggests, it is made from the redundant bits generated from next.
Define an encoding function mu which maps $\mathbf{u}_{i}\in{\{0,1\}}^{m}$ to a codeword $\textsc{mu}(\mathbf{u}_{i})\in{\{0,1\}}^{m+{\lceil\log m\rceil}+4}$ of a mutually-uncorrelated code $\mathcal{C}_{\text{MU}}$. The information region is hereby defined as
$\textsc{mu}(\mathbf{u}_{t})\circ\cdots\circ\textsc{mu}(\mathbf{u}_{l-1}).$
For $i\in[0,t-1]$, define the $i$-th redundancy packet
$\mathbf{d}_{i}=\mathbf{r}_{4i}\circ\mathbf{r}_{4i+1}\circ\mathbf{r}_{4i+2}\circ\mathbf{r}_{4i+3}$
as the concatenation of four redundancy strings. Then, define an encoding function rll which maps $\mathbf{d}_{i}$ to a binary sequence $\textsc{rll}(\mathbf{d}_{i})\in{\{0,1\}}^{4m+11}$ that is free of zero runs of length ${\lceil\log m\rceil}+1$ or more, and introduces $7$ bits of redundancy. The encoding rll gives rise to the following lemma which is easy to prove.
Lemma 1.
[…]
The redundancy region is hereby defined as
$\textsc{mu}(\mathbf{u}_{0})\circ\textsc{rll}(\mathbf{d}_{0})\circ\cdots\circ\textsc{mu}(\mathbf{u}_{t-1})\circ\textsc{rll}(\mathbf{d}_{t-1}).$
Finally, the codeword $\mathbf{c}$ is the two regions combined (line 12).
$\mathbf{c}$ $=\textsc{mu}(\mathbf{u}_{0})\circ\textsc{rll}(\mathbf{d}_{0})\circ\cdots\circ\textsc{mu}(\mathbf{u}_{t-1})\circ\textsc{rll}(\mathbf{d}_{t-1})$
$\phantom{=}\circ\textsc{mu}(\mathbf{u}_{t})\circ\cdots\circ\textsc{mu}(\mathbf{u}_{l-1}).$
Algorithm 2 Decode
1:Input: A multiset FRAGMENTS of at most $t+1$ (unordered) fragments of a codeword $\mathbf{c}\in\mathcal{C}$.
2:Output: The information word $\mathbf{w}$ such that Algorithm 1 with input $\mathbf{w}$ yields $\mathbf{c}$.
3:Let $\mathbf{r}_{i}\leftarrow\texttt{empty}$ for all $i\in[0,4t-1]$.
4:for all codeword $\textsc{mu}(\mathbf{u}_{i})\in\mathcal{C}_{\text{MU}}$ in the fragments in FRAGMENTS and $\mathbf{u}_{i}=\mathbf{p}^{(i)}$ for some integer $i<t$ do
5:     Let $\mathbf{m}_{i}$ be the $114m+$ 11 bits after $\textsc{mu}(\mathbf{u}_{i})$, and $\mathbf{d}_{i}\leftarrow\textsc{de-rll}(\mathbf{m}_{i})$.
6:     $\mathbf{r}_{4i}\leftarrow\mathbf{d}_{i}[0,m],\mathbf{r}_{4i+1}\leftarrow\mathbf{d}_{i}[m+1,2m+1],\mathbf{r}_{4i+2}\leftarrow\mathbf{d}_{i}[2m+2,3m+2],\mathbf{r}_{4i+3}\leftarrow\mathbf{d}_{i}[3m+3,4m+3]$
7:Let approxNext be a key-value store such that $\texttt{approxNext}[\mathbf{s}]=\mathbf{s}$ for all $\mathbf{s}\in{\{0,1\}}^{m}$.
8:for all fragment $\mathbf{f}=\mathbf{f}_{\text{start}}\circ\textsc{mu}(\mathbf{u}_{a})\circ\cdots\circ\textsc{mu}(\mathbf{u}_{a+s})\circ\mathbf{f}_{\text{end}}\in\texttt{FRAGMENTS}$ do
9:     for all $r\in[s-1]$ do $\texttt{approxNext}[\mathbf{u}_{a+r}]=\mathbf{u}_{a+r+1}$.
[…]
11:Let $\texttt{dStrings}=(\mathbf{u}_{1},\cdots,\mathbf{u}_{l})$ be an array of such that $\texttt{next}(\mathbf{u}_{i})=\mathbf{u}_{i+1}$.
[…]
13:return $\mathbf{u}[t\cdot m:]$
5.3 Decoding
Algorithm 2 provides a procedure for extracting the information word $\mathbf{w}$ from at most $t+1$ fragments of the respective codeword $\mathbf{c}$. The crux of this decoding procedure is to reconstruct the key-value store next (see Section 5.2), and based on which the information word $\mathbf{w}$ can be obtained. Specifically, the decoding algorithm creates a key-value store approxNext, which is slightly different from next, using the information which appears in the fragments. Alongside the correctly identified redundancy strings, approxNext goes through a Reed-Solomon decoding process and is corrected to next. Having the correct next in hand, the correct dStrings (Eq. (2)) can be found as described earlier. Then, dStrings is fed into d-decode (Alg. 4), which is the inverse process of d-encode (Alg. 3) to produce $\mathbf{u}$ (Eq. (1)), whose suffix is the information word $\mathbf{w}$.
The decoding starts by distinguishing and decoding the discernible codewords of $\mathcal{C}_{\textsc{mu}}$ from the fragments. Let $\textsc{mu}(\mathbf{u}_{i})$ be a discernible codewords in $\mathcal{C}_{\text{MU}}$ which fully reside within one fragment, where $\mathbf{u}_{i}$ is its respective decoding. If $\mathbf{u}_{i}=\mathbf{p}^{(i)}$ for some integer $i<t$, it means that $\mathbf{u}_{i}$ is a marker, and then the $(4m+11)$ bits after it is a redundancy packet (line 4). As such, the decoder passes the $(4m+11)$ bits after (if residing in the fragment) to an RLL-decoder and obtains four redundancy strings $\mathbf{r}_{4i},\mathbf{r}_{4i+1},\mathbf{r}_{4i+2},\mathbf{r}_{4i+3}$ (line 56). The process stops once no more such codewords can be found, and we have the following lemma.
Lemma 2.
Let $t_{2}$ be the number of breaks that fall in the redundancy region of the codeword. Then, the decoding algorithm is guaranteed to obtain $4(t-t_{2})$ redundancy strings.
Proof.
The decoding algorithm may fail to obtain a redundancy packet due to the occurrence of a cut either in the packet itself, or in the preceding marker. Therefore, it is guaranteed to obtain $t-t_{2}$ redundancy packets, and hence $4(t-t_{2})$ redundancy strings. ∎
The other discernible codewords of $\mathcal{C}_{\textsc{mu}}$, i.e., those encoded from non-markers ($\mathbf{u}_{i}$’s for $i\geq t$), are used to construct a key-value store approxNext. Initially, $\texttt{approxNext}(\mathbf{s})=\mathbf{s}$ for every key $\mathbf{s}\in{\{0,1\}}^{m}$ (line 7). For each fragment $\mathbf{f}$, let
$\mathbf{f}=\mathbf{f}_{\text{start}}\circ\textsc{mu}(\mathbf{u}_{a})\circ\cdots\circ\textsc{mu}(\mathbf{u}_{a+b})\circ\mathbf{f}_{\text{end}},$
where $\mathbf{f}_{\text{start}}$ and $\mathbf{f}_{\text{end}}$ are (possibly empty) prefix and suffix of $\mathbf{f}$ with no discernible codeword from $\mathcal{C}_{\text{MU}}$ that is encoded from a non-marker.
Then, in line 9, the decoder updates
$\texttt{approxNext}[\mathbf{u}_{a+c}]=\mathbf{u}_{a+c+1},$
for every $c\in[b-1]$. Note that approxNext is identical to next up to a bounded number of errors.
Lemma 3.
Let $t_{1}$ be the number of breaks that occur in the information region of the codeword. Then, approxNext differs from next at no more than $2t_{1}$ positions.
Proof.
The decoding algorithm may fail to identify a non-marker (i.e., $\mathbf{u}_{i}$ for $i\geq t$) in the information region if a break occurs inside it. Hence, the algorithm fails to identify at most $t_{1}$ non-markers. Notice that, failing to capture a non-marker $\mathbf{u}_{a}$ affects at most two key-value pairs of next, i.e., $\texttt{next}[\mathbf{u}_{a-1}]$ and $\texttt{next}[\mathbf{u}_{a}]$, and hence next and approxNext differ in at most $2t_{1}$ positions. ∎
The decoding algorithm proceeds to correct the constructed approxNext to next, i.e., the key-value store generated in Algorithm 1 from $\mathbf{w}$, using the collected redundancy strings and a standard Reed-Solomon decoder (line 10). Note that the missing redundancy strings are erasures (line 3), since their locations are known. The success of the decoding process is guaranteed in the following theorem.
Theorem 1.
Line 10 outputs the correct key-value store next.
Proof.
Notice that in line 10, a codeword is constructed by coalescing the elements in next with redundancy strings $\mathbf{r}_{0},\cdots,\mathbf{r}_{t-1}$, and fed into a Reed-Solomon decoder. By lemma 2, the decoder obtains at least $4(t-t_{2})$ redundancy strings, and as a result, the constructed codeword contains less than $4t_{2}$ erasures.
Meanwhile, Lemma 3 implies that the number of key-value pairs in which next and approxNext differ is bounded by $2t_{1}$. Hence, the systematic part of the constructed codeword has at most $2t_{1}$ errors. Recall that the encoding process generated $4t$ redundancy strings (symbols in $\mathbb{F}_{2}^{m+1}$), and hence the decoding in line 10 is guaranteed to be successful since
$2t_{1}\cdot 2+4\cdot t_{2}\leq 4(t_{1}+t_{2})\leq 4t,$
where the last transition follows since the actual number of breaks $t_{1}+t_{2}$ is at most the security parameter $t$. The proof is concluded since a $(2^{m}+4t,2^{m})$ Reed-Solomon code can simultaneously correct $a$ errors and $b$ erasures as long as $2a+b\leq 4t$. ∎
Finally, the array dStrings is obtained from next (line 11), and the function d-decode is employed to recover $\mathbf{u}$. Recall that $\mathbf{u}=\mathbf{p}^{(0)}\circ\cdots\circ\mathbf{p}^{(t-1)}\circ\mathbf{w}$, and hence the decoding procedure ends concludes by returning the $l\cdot m$ rightmost bits of $\mathbf{u}$ (line 13).
6 Implementation
We implemented a prototype of SIDE and conducted experiments to validate its applicability. First, the information bits $\mathbf{w}$ are encoded to a BRC codeword as described in Section 5.2. Then, we employ a Creality Ender 3 3D printer (Figure. 5(a)) to print objects with an embedded codeword (Figure 3(a)). The printer is operated by the Kipper firmware under the protection of OP-TEE. To read the information, a Leica S9D microscope is utilized to extract bits from the fragments of the printed object (Figure 3(b)). Note that reading externally using a visual device is not strictly necessary, as explained in the sequel. Finally, the bit strings that were recovered from the fragments go through the decoding process described in Section 5.3, and the original information is recovered.
6.1 BRC Encoding and Decoding
We implemented the BRC encoding and decoding procedures, whose pseudo-code is given in Algorithm 1 and Algoritm 2, respectively. The encoding part is written in C++ for the ease of using TEE. The implementation details and the use of TEE will be described in the future version of this paper. The decoding part is implemented in Python.
Figure 4: (a) A cuboid with an embedded BRC codeword. The codeword creates grooves on the outside of the object that are visible to the naked eye. For simplicity, we read the codeword using a microscope which detects light reflected from the top of each layer (Fig. 7). In cases where the outside grooves are not visible (e.g., scraping, filing, or coating), one can use a CT scanner to detect layer widths. (b) A cuboid broken by hand to $3$ fragments. Indeed, breaks occur naturally along cross-sections of the printing axis.
6.2 Information Embedding and Extraction
We argue that any practical bit embedding method should satisfy the following three properties.
  1. 1.
    Readability: The embedded information is readable within reasonable costs.
  2. 2.
    Robustness: The embedded information cannot be easily destroyed.
  3. 3.
    Obliviousness: The embedding method is oblivious to both the information bits and the geometry of the printed object.
The readability requirement implies economical reading of bits from the broken object with respect to a given use case. The robustness requirement states that the embedded information must be sufficiently robust against adversarial tempering, within a certain bound. For example, the surface of the printed object may be easily abraded and as a result, embedding information on the surface is not robust.
The obliviousness requirements decouples the embedding method both from the information to be embedded and from the 3D models. Examples which violate the obliviousness requirement include representing $1$’s by thick layers and representing 00’s by thin layers; this will cause codewords with different number of 1’s to have embeddings of different lengths. Moreover, embedding bits by alerting the geometry of the printed object is also considered inappropriate, as it may significantly affect its functionality. Note that for any embedding method, imposing some requirement on the dimension of the 3D model is inevitable (e.g., that its size is sufficient for embedding a certain number of bits), and we do not consider this as a violation of the obliviousness requirement.
ptptptptptpt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
ptptpt
Figure 5: Demonstration of embedding $01101001$ by varying layer width, in which each 00 is represented by three layers of $0.12$ millimeters, and each $1$ is represented by two layers of $0.18$ millimeters. The length required for embedding one bit is irrelevant to its value.
In our implementation we embed bits by varying layer width of the printed object, with a slight modification. As shown in Figure 5, a 00 is represented by three layers of $0.12$ millimeters, and a $1$ is represented by two layers of $0.18$ millimeters. These values are chosen to maximize information density, i.e., the number of bits that can be embedded per unit length, with respect to the minimum layer height ($0.12$ millimeters) offered by the Ender 3 printer used in the experiments. As the heights add up to $0.36$ millimeters in both cases, this method avoids the situation of height discrepancy mentioned in the previous paragraph, and satisfies the obliviousness requirement. As for the robustness requirement, we note that layer height is an intrinsic property of the printed object, and cannot be easily altered, nor it is affected by breaking or scraping the object.
The reading of information from fragments of the printed object is done by a Leica S9D microscope, which provides a high resolution that captures their rugged surface. It is worth noting that the reading of bits does not require the surface of the fragment to be free of adversarial tempering, as less-economical solutions, e.g., using an industrial CT scanner, can be used to infer the layer width via tomographic analysis.
Figure 6: (a) The creality Ender 3 3D printer used in prototyping and experiments. (b) The fragment moves with the motorized slider, and the Leica S9D microscope takes a series of images of different parts of it.
Figure 7: The reading of a fragment (cropped due to space limit), in which every bright lines is the center of a layer. The pixel values (in gray scale) are accumulated and normalized for every column and displayed in red. The peaks are marked by yellow dots, which reveal the distance between every two consecutive bright lines.
Due to the limitation of field of view, the microscope is unable to capture the entire fragment all at once. To address this, the fragment to be examined is fixed on a motorized rail slider (see Figure 5(b)). During the information extraction process, the fragment slides over the field of view of the microscope, which takes series of pictures in the meantime. Note that the pictures are taken so that every two consecutive pictures overlap. This feature allows us to fuse them together and obtain a picture of the entire fragment using ImageJ with the stitching plugin [39].
In the resulting stitched image, shown in Figure 7, the bright lines are the center of each layer. To automatically read the bits from this image, we implemented a program that reads the distance between every two adjacent lines, and places them in an array. Recall that there are two types of layers, the thin ($0.12$ mm) and the thick ($0.18$ mm). Hence, the elements in the array can be placed in three categories, i.e., two adjacent thin layers (thin-thin), a thin layer followed by a thick one or vice versa (thin-thick), and two adjacent thick layers (thick-thick). The program then applies $K$-means clustering with $K=3$ in order to differentiate the three categories.
Next, the program examine the clustering results and corrects potential errors resulting from the lighting environment during microscope reading or the image distortion in stitching. The embedding method (Figure 5) imposes certain patterns, which allow such correction. For example, a thin-thin distance must appear adjacent to another thin-thin distance, and in between thin-thin and thick-thick there must be a thin-thick.
For each layer, the program determines whether it is a thick one or a thin one by the distances from its adjacent lines, e.g., a thin line may have a thick-thick distance on its left and a thick-thin distance on its right. This concludes the reading of bits from fragments.
# of breaks information length codeword length information rate object length information density
1 98 bits 243 bits 0.403 87.48 mm 1.089 bits/mm
2 120 bits 353 bits 0.340 127.08 mm 0.944 bits/mm
Table I: The evaluation is conducted on two groups of five sample prints. In the first group, we embed 98 bits of information with a $1$-break code, which results in codewords of 243 bits and requires the object to be at least $\approx 87.48$mm along the printing direction. Sample prints in the second group are resilient to $2$ breaks, and the $2$-break code in use has information length $k=120$ and codeword length $n=353$. At least $\approx 127.08$mm along the print direction are required for the object in this case.
7 Evaluation
The evaluation of our prototype focuses on the information density (number of bits per unit distance, which is directly affected by the codeword length $n$ of BRC) and the recoverability of data from broken objects. The former measures the amount of information that can be embedded in an given object, and the latter provides the secure information extraction feature of SIDE.
On the theoretical side, we analyzed the growth of the codeword length $n$ as a function of $k$ and $t$, for $k\in\{64,96,128,160\}$ and $t\in\{1,2,\ldots,9\}$ (Figure 8). On the engineering side, we conduct experiments on two groups of samples, which differ in the number of embedded bits (reflected by $k$) and the number of breaks that can be tolerated (reflected by $t$).
For the first group, we printed five sample cuboids of dimension $6$mm×$\times 20$mm×$\times 90$mm, and implemented a $1$-break code with $k=98$ and $n=243$, which maps an information word of $98$ bits to a $243$ bit codeword. For each cuboid, we prepared a distinct information string and encoded it to a codeword of the aforementioned $1$-break code; the codeword is embedded in the cuboid and is decodable even if broken at most once at any arbitrary location. For the second group, we develop a $2$-break code with $k=120$ and $n=353$. The sample cuboids in this group has a dimension of $8$mm×$\times 20$mm×$\times 130$mm. The cuboids in this group are resilient to two breaks in any arbitrary location. Table I details the parameters of both groups.
For every cuboid, we broke it apart by hand into $t+1$ parts, where $t$ is the number of breaks to which it is resilient. We examined the fragments and read bits (see Section 6.2). We assumed that the orientation of the fragments (i.e., the correct $z$ direction) is known to the decoder; we postulate that closer examination of individual layers can alleviate this assumption due to the impact of gravity on the molten thermoplastic, and leave this to future work. The bit strings from the fragments were fed into the decoding algorithm, and the decoding results are compared against the original encoded data. The information extraction succeeded in all ten cases.
00111122223333444455556666777788889999002002002002004004004004006006006006008008008008001,00010001{,}0001 , 000number of breaks $t$codeword length $n$Demonstration of $t,k,n$$k=64$$k=96$$k=128$$k=160$
Figure 8: Demonstration of $t,k,n$
8 Conclusion
In this paper, we present SIDE, which aims to provide secure information embedding and extraction in forensic 3D fingerprinting against the malicious tempering of both 3D printers and the printed objects. SIDE is built upon hardware isolation mechanisms and coding-theoretic techniques. we develop break-resilient code, which allows for information extraction from the printed object even if it is broken into pieces. We leverage trusted execution environment to protect the integrity of the printing process, which guarantees the embedding of the break-resilient codeword. We implement a prototype of SIDE and validated our design.
Finally, we propose future studies in order to advance secure 3D fingerprinting technology further. On the engineering side, we propose to study the integration of trusted execution environments in commodity printers, and the trusted procurement of bits (e.g., timestamp or geolocation) for the sake of embedding. Another engineering challenge is coming up with alternative secure methods for bit embedding in order to increase information density. On the theoretical side we propose to devise practical coding methods which operate closer to the information-theoretic limits of this model [26], as well as more sophisticated methods for reconstructing the information from only a subset of fragments, rather than all of them.
9 Acknowledgment
This research was supported by the National Science Foundation under Grant CNS-2223032.